# LLM-Forest: Ensemble Learning of LLMs with Graph-Augmented Prompts for Data Imputation

### v3

@@ Abstract
Missing data imputation is a critical challenge in various domains, such as healthcare and finance, where data completeness is vital for accurate analysis. Large language models (LLMs), trained on vast corpora, have shown strong potential in data generation, making them a promising tool for data imputation. However, challenges persist in designing effective prompts for a finetuning-free process and in mitigating the risk of LLM hallucinations. To address these issues, we propose a novel framework, LLM-Forest, which introduces a "forest" of few-shot learning LLM "trees" with confidence-based weighted voting, inspired by the ensemble learning (Random Forest). This framework is established on a new concept of bipartite information graphs to identify high-quality relevant neighboring entries with both feature and value granularity. Extensive experiments on 9 real-world datasets demonstrate the effectiveness and efficiency of LLM-Forest.
@@ 1 Introduction
Handling missing data is fundamental across various domains, including datasets such as clinical reports, financial records, product characteristics, and survey data in a tabular format Sterne et al. (2009); Shwartz-Ziv and Armon (2022); Hernandez et al. (2022). However, these datasets often suffer from missing values, which pose significant challenges for critical decision-making for studies that rely on complete data for accurate predictions and analyses. To address this issue, feature imputation Brick and Kalton (1996); Troyanskaya et al. (2001); Jazayeri et al. (2020); Bernardini et al. (2023)—the process of estimating missing values based on observed data—plays a crucial role in maintaining data integrity and ensuring meaningful analysis.
@@
Common approaches for feature imputation can be broadly categorized into two groups. (1)Statistical Methods: Techniques such as mean or mode imputation and MICE Little and Rubin (2019) are simple and widely used due to their ease of implementation. While these methods are straightforward, they can introduce bias. (2)Deep Learning-Based Methods: Models like GAIN Yoon et al. (2018), DIFFIMPUTE Mattei and Frellsen (2019), and GRAPE You et al. (2020) aim to model complex data distributions and dependencies leveraging advanced neural network architectures to capture non-linear relationships within the data. However, they typically require large volumes of data and can be challenging to train, particularly in scenarios like healthcare where data is often limited, imbalanced, or subject to strict privacy regulations.
@@
Recently, Large Language Models (LLMs) Brown (2020); Devlin (2018); Beltagy et al. (2019); Agrawal et al. (2023) have demonstrated strong potential in handling complex tasks by leveraging external knowledge and contextual learning, offering a promising solution for enhancing performance on tabular data tasks Gong et al. (2020); Yin et al. (2020); Abraham et al. (2022); Thirunavukarasu et al. (2023). Anam Nazir et al. (2023) began exploring the adoption of ChatGPT Brown (2020) as a data imputer by formulating text questions and prompting ChatGPT to respond with imputed values for each missing cell. Similarly, CLAIM Hayat and Hasan (2024) introduced an LLM approach for generating context-specific descriptions of missing data. However, both methods require fine-tuning processes, which can be time-consuming and resource-intensive, limiting their scalability across different datasets. In-context learning Wei et al. (2022); Min et al. (2022); Gruver et al. (2024) has excelled in downstream tasks like tabular classification (e.g., CancerGPT Li et al. (2024)) and reasoning (e.g., CHAIN-OF-TABLE Wang et al. (2024)), but prompt-based data imputation remains largely unexplored.
@@
In this paper, we explore fine-tuning-free approaches to harness the power of LLMs for data imputation. We aim to overcome two primary challenges: (1) extracting high-quality information from observed data to align with the preferences of the target entries, and (2) mitigating the hallucinations inherent in current LLMs.
@@
To address these challenges, we propose a novel framework, LLM-Forest. It consists of multiple trees utilizing few-shot in-context learning with LLMs, inspired by ensemble learning (Random Forest), reducing the chance of hallucination compared to a single LLM. Each tree in the forest represents an LLM with a particular few-shot prompt learning process. To construct high-quality but diverse prompts, we design a graph-based retrieval algorithm to explore highly relevant neighboring entries in preparation for constructing LLM-Forest. This graph algorithm is based on Information-Theoretic Bipartite Graphs to quantify relevance at both the feature-level and value-level granularity. Meanwhile, it reduces the computational complexity in the graph construction process and allows for the subsequent merging process to be easily implemented in parallel. This makes it scalable for very large datasets. Finally, we aggregate the outputs of the LLM trees by weighting each tree’s outcome according to its confidence level, ensuring more robust and reliable results. In summary, our main contributions are :
@@
  • Information-based Graph retrieval for LLM: We integrate information theory into the construction of bipartite graphs to accommodate different distributions and value-level granularity, along with an efficient merging process. The random-walk-based retrieval adds randomness (exploration) to each LLM tree, ensuring that the trees are less correlated with each other.
@@ 2 Notations and Problem Definitions
Let $[n]=\{1,2,\dots,n\}$ and define $N=[n]$. Suppose the matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents a tabular dataset with $n$ entries and $d$-dimensional features. The matrix $\mathbf{X}=[A_{1},A_{2},\dots,A_{n}]^{\top}$, where each entry $A_{i}\in\mathbb{R}^{d}$, $i\in[n]$, is a $d$-dimensional vector. For example, in a diabetes dataset, $A_{i}^{\top}$ could represent the feature set of the patient $i$. Let $A_{ij}$, $j\in[d]$, denote the $j$-th feature value of the $i$-th entry $A_{i}^{\top}$. For convenience, we use $\{C_{1},C_{2},\dots,C_{d}\}$ to represent the $d$-dimensional features (columns) of $\mathbf{X}$, where $C_{ji}\in\mathbb{R}^{n}$, $j\in[d],i\in N$, denotes the $i$-th entry of the $j$-th feature. Let $R_{j}$ represent the set of distinct values in feature $C_{j}$, which can be viewed as the collection of attribute values for all patients in the dataset. Additionally, define a matrix $\mathbf{M}\in\mathbb{R}^{n\times d}$ to represent the missingness pattern of $\mathbf{X}$, where a cell has the value 1 if the corresponding entry in $\mathbf{X}$ is missing, and 0 otherwise.
@@
Figure 1: Overview of LLM-Forest. LLM-Forest  consists of four key steps: (1) Table to Bipartite Information Graphs: This step converts tabular data into bipartite information graphs, capturing the relationships between dimensional features and their corresponding entries. (2) Merge Bipartite Graphs: The bipartite graphs are then hierarchically merged, aggregating shared information across different features. (3) Random Walk: Random walks are performed on the merged bipartite graphs to identify correlated and diverse neighbor nodes for each entry (node), preparing for the construction of the LLM-tree. (4) Construct LLM-tree to Forest: Based on the identified neighbors, a tailored few-shot prompt is constructed for each LLM, forming a forest with confidence-weighted voting.
@@ 3 Proposed Method: LLM-Forest
In this section, we introduce the proposed framework, LLM-Forest, which consists of four key steps, described by Figure 1. We detailed each procedure below.
@@ 3.1 Table to Bipartite Information Graphs
Given an entry $A_{i}$, a dimension $C_{j}$, and a value $a_{k}^{j}\in R_{j}$, suppose the probability of the event that $A_{ij}=a_{k}^{j}$ is $p_{i}(a_{k}^{j})$. According to information theory Ash (2012), the self-information of the event that entry $A_{i}$ takes the value $a_{k}^{j}$ in dimension $j$ is defined as:
@@
Determine $p_{i}(a^{k}_{j})$. The next critical step is to determine $p_{i}(a_{k}^{j})$. Below, we provide three formulas based on the different types of features.
@@
3.2 Merge Bipartite Graphs
We use a bottom-up hierarchical process to merge the $d$ bipartite graphs into compressed bipartite graphs. In each step, two bipartite graphs are merged into one, while accumulating the shared information of entries across the two dimensional features. Given two bipartite graphs $G_{j}=(N\cup R_{j},E_{j})$ and $G_{j^{\prime}}=(N\cup R_{j^{\prime}},E_{j^{\prime}})$, where $j,j^{\prime}\in[d]$, we merge the right sides $R_{j}$ and $R_{j^{\prime}}$, as both graphs share the same left side $N$.
@@ 3.2 Merge Bipartite Graphs
This rule ensures that the information shared by entry $i$ is accumulated across both dimensions. Note that $w_{i}(a_{k}^{j})=0$ if the edge $e_{i}(a_{k}^{j})$ does not exist. By merging nodes in descending order of the Jaccard coefficient $S$, a new bipartite graph $G_{\hat{j}}=(N\cup R_{\hat{j}},E_{\hat{j}})$ is formed.
@@
3.3 Random Walk
@@ 3.4 Construct LLM-tree to Forest
Promept design. To bridge the gap between the structured format of tabular data and the natural language processing capabilities of LLMs, our methodology incorporates a table transformation module. This module utilizes a simple template to convert structured tabular data into a natural language format. For each entry (representing an individual sample), we extract the data across various features, such as age, gender, smoking status, and more from the structured dataset. These feature-value pairs are then formatted into a prompt using the template: for each feature, if a value is present, it is included as "feature: value;" and we add a statement as "the sample has missing features: missing feature 1, missing feature 2, …." at the end for all the missing values.
@@
Confidence-weighted Voting. The most straightforward method for determining the ensemble output of the LLM Forest is majority voting. However, we propose an enhanced strategy by weighting each LLM-tree’s vote according to its confidence level. Current LLMs can provide a confidence level associated with their inferences for specific tasks. For example, GPT4 assigns "High", "Medium" or "Low" confidence levels to its inference output. A higher confidence reflects a stronger alignment between the model’s output and its internal logic and prior knowledge. Therefore, we introduce a criterion where higher confidence results in a greater voting weight. For instance, LLM trees with "High" confidence receive a voting weight of 1.0, while those with "Medium" confidence are assigned a weight of 0.6, "Low" with 0.3. The final imputed value is determined by selecting the value with the largest cumulative weighted votes.
@@ 4.1 Exepriment setup
We evaluate the imputation performance of the proposed method using 9 datasets from different domains: NPHA nat (2017), Gliomas Tasci and Zhuge (2022), Cancer Borzooei and Tarokhian (2023), Credit-g Hofmann (1994), Concrete Yeh (1998), Yacht Gerritsma and Versluis (1981) and Wine Cortez and Reis (2009) datasets from the UCI Machine Learning Repository and Diabetes and Housing Harrison Jr and Rubinfeld (1978) dataset from Kaggle. The detailed information datasets, the preprocessing procedures, the construction of the training and test sets, and the evaluation metrics are elaborated in Appendix A.2.
We compare LLM-Forest with the widely used imputation baselines including both statistic and deep learning-based methods as described in Appendix A.2. We also introduce zero-shot LLM (without similar samples provided): LLM-zero and adapt Chain-of-Thought Wei et al. (2022) framework by providing example questions alongside corresponding imputation results and explanations for LLMs for comparison. The method proposed in Nazir et al. (2023) generating a separate question for each missing value can be considered a variant of the LLM-zero.
Main experiments are conducted with GPT-4o (referred to as GPT-4 in the following discussion) and we also provided the results with Claude-3.5 and open-source model Mixtral-8×22B-v0.1 in the Appendix A.3. The experiments are conducted 3 times to ensure reliability. The performance improvements are statistically significant with p-values less than 0.05 when compared to baseline methods. The hyperparameter settings for LLM-Forest and baseline methods are provided in Appendix A.2.
@@ 4.2 Main Results & Discussion
Imputation Performance. We begin by evaluating the imputation accuracy of our proposed methods, as presented in Table 1. LLM-Forest achieves the highest imputation performance across all datasets when compared to traditional imputation techniques and LLM-based in-context learning baselines. From the main experiments, we derive the following insights:
(i) LLMs exhibit robust performance across datasets with varying feature distributions. For instance, In Gliomas, over 70% of features are highly skewed, with over 90% of values in one category. Despite this, LLM-Forest still outperforms statistical baselines. (ii) LLMs are effective for imputing continuous datasets. As shown in Table 1, LLM-Forest achieves the lowest MAE across all continuous datasets. These results demonstrate the capability of LLMs to generate accurate values when faced with a wide range of possible outcomes in continuous imputation tasks. (ii) A single LLM (LLM-Tree), when provided with an appropriately designed prompt, is capable of delivering accurate imputations. Across all datasets, the LLM-Tree achieves performance within less than 1% of the best method, consistently maintaining a high level of accuracy. (iii) LLMs’ own knowledge is limited, and they need to learn from informative contexts to perform well in this task. In the zero-shot LLM setting, without informative neighbors, LLM-Zero leaves 9.27% of Gliomas and 9.52% of Diabetes cells unimputed. It also produces uncertain or unreasonable outputs, like "approximately" or "18000" for Age. (iv) Forming an LLM forest enhances the robustness of imputation results. Compared to a single LLM, using just three trees can lead to performance improvements across all datasets. This highlights the effectiveness of ensemble learning in improving imputation robustness and accuracy of LLMs.
Performance on Downstream tasks. We evaluate the downstream classification task performance using a logistic regression model Cox (1958) on five datasets. As shown in Figure 2, LLM-Forest consistently achieves the highest performance across 4 evaluated datasets and second-best on Gliomas. This superior performance can be attributed to the LLMs’ ability to identify and leverage relevant patterns during imputation, as well as its capacity to effectively apply its extensive base knowledge.
@@
$ACC(\uparrow)$ $ACC(\uparrow)$ $ACC(\uparrow)$ $ACC(\uparrow)$ $ACC(\uparrow)$ $MAE(\downarrow)$ $MAE(\downarrow)$ $MAE(\downarrow)$ $MAE(\downarrow)$

### v4

@@ Abstract
Missing data imputation is a critical challenge in various domains, such as healthcare and finance, where data completeness is vital for accurate analysis. Large language models (LLMs), trained on vast corpora, have shown strong potential in data generation, making them a promising tool for data imputation. However, challenges persist in designing effective prompts for a finetuning-free process and in mitigating biases and uncertainty in LLM outputs. To address these issues, we propose a novel framework, LLM-Forest, which introduces a "forest" of few-shot prompt learning LLM "trees" with their outputs aggregated via confidence-based weighted voting based on LLM self-assessment, inspired by the ensemble learning (Random Forest). This framework is established on a new concept of bipartite information graphs to identify high-quality relevant neighboring entries with both feature and value granularity. Extensive experiments on 9 real-world datasets demonstrate the effectiveness and efficiency of LLM-Forest. The implementation is available at https://github.com/Xinrui17/LLM-Forest
@@ 1 Introduction
Handling missing data is a fundamental challenge across various domains, including datasets such as clinical reports, financial records, product characteristics, and survey data in a tabular format Sterne et al. (2009); Sankar et al. (2021); Shwartz-Ziv and Armon (2022); Hernandez et al. (2022); Wang et al. (2024a). However, these datasets frequently suffer from missing values, which can severely impact the reliability of critical decision-making processes that depend on complete and accurate data. To address this issue, feature imputation Brick and Kalton (1996); Troyanskaya et al. (2001); Jazayeri et al. (2020); Bernardini et al. (2023), the process of estimating missing values based on observed data, plays a crucial role in maintaining data integrity and ensuring meaningful analysis.
Figure 1: Task-agnostic framework of LLM-Forest: Ensemble of LLM outputs based on generated diverse prompts for a specific task.
@@
Common approaches for feature imputation can be broadly categorized into two groups. (1) Statistical Methods: Techniques such as mean or mode imputation and MICE Little and Rubin (2019) are simple and widely used due to their ease of implementation. While these methods are straightforward, they can introduce bias. (2) Deep Learning-Based Methods: Models like GAIN Yoon et al. (2018), DIFFIMPUTE Mattei and Frellsen (2019), and GRAPE You et al. (2020) aim to model complex data distributions and dependencies leveraging advanced neural network architectures to capture non-linear relationships within the data. However, they typically require large volumes of data and can be challenging to train, particularly in scenarios like healthcare where data are often limited, imbalanced, or subject to strict privacy regulations.
@@
Recently, Large Language Models (LLMs) Brown (2020); Devlin (2018); Beltagy et al. (2019); Agrawal et al. (2023); Zou et al. (2025a); Huang et al. (2025); Li et al. (2025); Ai et al. (2025) have demonstrated strong potential in handling complex tasks by using external knowledge and contextual learning, offering a promising solution for enhancing performance on tabular data tasks Gong et al. (2020); Yin et al. (2020); Abraham et al. (2022); Thirunavukarasu et al. (2023). Anam Nazir et al. (2023) began exploring the adoption of ChatGPT Brown (2020) as a data imputer by formulating text questions and prompting ChatGPT to respond with imputed values for each missing cell. Similarly, CLAIM Hayat and Hasan (2024) introduced an LLM-based approach to generate text-specific descriptions of missing data. However, both methods require fine-tuning processes, which can be time-consuming and resource-intensive, limiting their scalability across different datasets. In contrast, in-context learning Wei et al. (2022); Min et al. (2022); Gruver et al. (2024); Li et al. (2024b) has excelled in tasks such as tabular classification (e.g., CancerGPT Li et al. (2024a)) and reasoning (e.g., CHAIN-OF-TABLE Wang et al. (2024b)). Yet the use of LLMs’ broad knowledge and reasoning ability for data imputation, without relying on task-specific fine-tuning, remains largely unexplored.
@@
In this paper, we explore fine-tuning-free approaches to harness the power of pre-trained LLMs for data imputation. We aim to overcome two primary challenges: (1) extracting high-quality information from observed data to align with the preferences of the target entries to effectively guide LLMs, and (2) reduce potential output bias and uncertainty in current LLMs.
To address these challenges, we propose a general and scalable framework, LLM-Forest. It consists of multiple LLM-based trees, each guided by a distinct few-shot prompt designed to capture different perspectives of the task. These prompts can be constructed using various strategies, such as LLM-based generation, graph search, or specific algorithms. Inspired by ensemble learning like Random Forest, LLM-Forest aggregates the predictions from all trees using confidence-based weighting, where each LLM estimates its own output confidence. As shown in Figure 1, this task-agnostic design supports a wide range of applications including question answering, text generation, and classification, offering improved robustness and reduced bias compared to using a single LLM.
@@
In this paper, we instantiate LLM-Forest for the task of tabular data imputation, where the goal is to fill in missing values based on observed entries. We demonstrate how the general design of LLM-Forest can be tailored to imputation through graph-based neighbor selection. Each tree in the forest represents an LLM with a particular few-shot prompt learning process. To construct high-quality yet diverse prompts, we design a graph-based retrieval algorithm to explore highly relevant neighboring entries in preparation for constructing LLM-Forest. This graph algorithm is based on Information-Theoretic Bipartite Graphs to quantify relevance at both the feature-level and value-level granularity. Meanwhile, it reduces the computational complexity in the graph construction process and allows for the subsequent merging process to be easily implemented in parallel. This makes it scalable for very large datasets. We then carefully design prompts to integrate the retrieved neighbors and inject domain-level guidance, maximizing the utility of pre-trained LLMs without any fine-tuning. Finally, we aggregate the outputs of the LLM trees by weighting each tree’s prediction according to its self-assessed confidence level, ensuring more robust and reliable results. In summary, our main contributions are :
@@
  • Information-based Graph retrieval for LLM: We integrate information theory into the construction of bipartite information graphs to accommodate different distributions and value-level granularity, along with an efficient merging process. Random walk-based retrieval adds randomness (exploration) to each LLM tree, ensuring that the trees are less correlated with each other.
@@
Figure 2: Overview of LLM-Forest for data imputation. LLM-Forest  consists of four key steps: (1) Bipartite Information Graph construction: converts tabular data into bipartite information graphs, capturing the relationships between data entries and the values of their associated dimensional features. (2) Merge Bipartite Graphs: Hierarchically merges the bipartite information graphs to aggregate shared information across features. (3) Graph Retrieval via Random Walk: Performed on the merged graphs to identify correlated and diverse neighbors for each entry. (4) Construct LLM-tree to Forest: Builds a tailored few-shot prompt using the identified neighbors for each LLM; Multiple LLMs form a “forest” whose outputs are aggregated via confidence-weighted voting to generate the final imputed values.
@@ 2 Notations and Problem Definitions
Let $n]=\{1,2,\dots,n\}$ and define $N=[n]$. Suppose the matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents a tabular dataset with $n$ entries and $d$-dimensional features. The matrix $\mathbf{X}=[A_{1},A_{2},\dots,A_{n}]^{\top}$, where each entry $A_{i}\in\mathbb{R}^{d}$, $i\in[n]$, is a $d$-dimensional vector. For example, in a diabetes dataset, $A_{i}^{\top}$ could represent the feature set of the patient $i$. Let $A_{ij}$, $j\in[d]$, denote the $j$-th feature value of the $i$-th entry $A_{i}^{\top}$. For convenience, we use $\{C_{1},C_{2},\dots,C_{d}\}$ to represent the $d$-dimensional features (columns) of $\mathbf{X}$, where $C_{ji}\in\mathbb{R}^{n}$, $j\in[d],i\in N$, denotes the $i$-th entry of the $j$-th feature. Let $R_{j}$ represent the set of distinct values in feature $C_{j}$, which can be viewed as the collection of attribute values for all patients in the dataset. Additionally, define a matrix $\mathbf{M}\in\mathbb{R}^{n\times d}$ to represent the missingness pattern of $\mathbf{X}$, where a cell has the value 1 if the corresponding entry in $\mathbf{X}$ is missing, and 0 otherwise.
@@ 3 Proposed Method: LLM-Forest
In this section, we present our proposed framework, LLM-Forest, which consists of four key components, as illustrated in Figure 2. The first three steps form the information-based graph retrieval module: (1) Table to Bipartite Information Graphs: This step converts tabular data into bipartite information graphs connecting dimensional features and their corresponding entries, where edges are weighted based on self-information to quantify the informativeness of feature values, enriching the representation of feature–entry relationships. (2) Hierarchical Bipartite Graph Merging: The bipartite graphs are then hierarchically merged, aggregating shared information across different features. (3) Graph Retrieval via Random Walk: Random walks are performed on the merged bipartite graphs to identify correlated and diverse neighbor nodes for each entry (node), preparing for the construction of the LLM-tree. Finally, (4) Construct LLM-tree to Forest: Based on the identified neighbors, a tailored few-shot prompt is constructed for each LLM, forming a forest with confidence-weighted voting. We detailed each procedure below.
@@ 3.1 Table to Bipartite Information Graphs
Given an entry $A_{i}$, a dimension $C_{j}$, and a value $a_{k}^{j}\in R_{j}$, suppose that the probability of the event $A_{ij}=a_{k}^{j}$ is $p_{i}(a_{k}^{j})$. Grounded in information theory Ash (2012), the self-information of the event that entry $A_{i}$ takes the value $a_{k}^{j}$ in dimension $j$ is defined as:
@@
Determine $p_{i}(a^{k}_{j})$. To compute the informativeness-based edge weight $w_{i}(a_{k}^{j})$, we estimate the probability $p_{i}(a_{k}^{j})$ based on the type of feature $C_{j}$. We consider three cases:
@@
3.2 Hierarchical Bipartite Graph Merging
We use a bottom-up hierarchical process to merge the $d$ bipartite graphs into compressed bipartite information graphs. In each step, two bipartite graphs are merged into one, while accumulating the shared information of entries across the two dimensional features. Given two bipartite graphs $G_{j}=(N\cup R_{j},E_{j})$ and $G_{j^{\prime}}=(N\cup R_{j^{\prime}},E_{j^{\prime}})$, where $j,j^{\prime}\in[d]$, we merge the right sides $R_{j}$ and $R_{j^{\prime}}$, as both graphs share the same left side $N$.
@@ 3.2 Hierarchical Bipartite Graph Merging
[…]
@@
3.3 Graph Retrieval via Random Walk
@@ 3.4 Construct LLM-tree to Forest
Promept design. To bridge the gap between the structured format of tabular data and the natural language processing capabilities of LLMs, our methodology incorporates a table transformation module. This module utilizes a simple template to convert structured tabular data into a natural language format. For each entry (representing an individual sample), we extract the data across various features, such as age, gender, smoking status, and more from the structured dataset. These feature-value pairs are then formatted into a sample using the template: for each feature, if a value is present, it is included as "feature: value;" and we add a statement as "the sample has missing features: missing feature 1, missing feature 2, …." at the end for all the missing values.
@@
Confidence-weighted Voting. The most straightforward method for determining the ensemble output of the LLM Forest is majority voting. However, leveraging the self-assessment ability, we propose an enhanced strategy by weighting each LLM-tree’s vote according to its confidence level. Current LLMs can provide a confidence level associated with their inferences for specific tasks. For example, GPT4 assigns "High", "Medium" or "Low" confidence levels to its inference output. A higher confidence reflects a stronger alignment between the model’s output and its internal logic and prior knowledge. Therefore, we introduce a criterion where higher confidence results in a greater voting weight. For instance, LLM trees with "High" confidence receive a voting weight of 1.0, while those with "Medium" confidence are assigned a weight of 0.6, "Low" with 0.3. The final imputed value is determined by selecting the value with the largest cumulative weighted votes.
@@ 4.1 Exepriment setup
We evaluate the imputation performance of the proposed method using 9 datasets from different domains: NPHA nat (2017), Gliomas Tasci and Zhuge (2022), Cancer Borzooei and Tarokhian (2023), Credit-g Hofmann (1994), Concrete Yeh (1998), Yacht Gerritsma and Versluis (1981) and Wine Cortez and Reis (2009) datasets from the UCI Machine Learning Repository and Diabetes and Housing Harrison Jr and Rubinfeld (1978) dataset from Kaggle. The detailed information on datasets, the preprocessing procedures, the construction of the training and test sets, and the evaluation metrics are elaborated in Appendix A.2.
We compare LLM-Forest with the baselines including both statistic and deep learning-based methods as described in Appendix A.2. We also introduce zero-shot LLM (without similar samples provided): LLM-zero and adapt Chain-of-Thought Wei et al. (2022) framework by providing example questions alongside corresponding imputation results and explanations for LLMs for comparison. The method proposed in Nazir et al. (2023) generating a separate question for each missing value can be considered a variant of the LLM-zero.
Main experiments are conducted with GPT-4o (referred to as GPT-4 in the following discussion) and we also provided the results with Claude-3.5 and open-source model Mixtral-8×22B-v0.1 in the Appendix A.3. The experiments are conducted 3 times to ensure reliability. The performance improvements are statistically significant with p-values less than 0.05 when compared to baselines. The hyperparameter settings for LLM-Forest and baseline methods are provided in Appendix A.2.
@@ 4.2 Main Results & Discussion
Imputation Performance. We begin by evaluating the imputation accuracy of our proposed methods, as presented in Table 1. LLM-Forest achieves the strong imputation performance across all datasets when compared to traditional imputation techniques and LLM-based in-context learning baselines. From the main experiments, we derive the following insights:
(i) LLMs exhibit strong performance for data imputation across datasets with varying feature distributions. For instance, In Gliomas, over 70% of features are highly skewed, with over 90% of values in one category. Despite this, LLM-Forest still outperforms statistical baselines. (ii) LLMs are effective for imputing continuous datasets. As shown in Table 1, LLM-Forest achieves the lowest MAE on most continuous datasets and ranks second on the remaining one. These results demonstrate the capability of LLMs to generate accurate values when faced with a wide range of possible outcomes in continuous imputation tasks. (ii) A single LLM (LLM-Tree), when provided with an appropriately designed prompt, is capable of delivering accurate imputations. Across all datasets, the LLM-Tree achieves performance within less than 1% of the best method, consistently maintaining a high level of accuracy. (iii) LLMs’ own knowledge is limited, and they need to learn from informative contexts to perform well in this task. In the zero-shot LLM setting, without informative neighbors, LLM-Zero leaves 9.27% of Gliomas and 9.52% of Diabetes cells unimputed. It also produces uncertain or unreasonable outputs, like "approximately" or "18000" for Age. (iv) Forming an LLM forest enhances the accuracy and robustness of imputation results. Compared to a single LLM, using just three trees can lead to performance improvements across all datasets. This highlights the effectiveness of ensemble learning in improving imputation robustness and accuracy of LLMs.
Performance on Downstream tasks. We evaluate the downstream classification task performance using a logistic regression model Cox (1958) on five datasets. As shown in Figure 3, LLM-Forest consistently achieves the highest performance across 4 datasets and second-best on Gliomas. This superior performance can be attributed to the LLMs’ ability to identify relevant patterns among features and entries during imputation, as well as its capacity to effectively apply its extensive base knowledge.
@@
[…]